complete combustion of 7.30 g of a hydrocarbon produced 22.5 g of CO2 and 10.4 g of H2O. What is the empirical formula for the hydrocarbon
1
Expert's answer
2020-09-09T04:36:44-0400

Solution

The key here is to realize that you're dealing with a hydrocarbon, that is, a compound that contains only carbon and hydrogen.

Notice that the products of this combustion reaction are carbon dioxide,CO2, and water, H2O.


This tells you that all the carbon that was initially a part of the hydrocarbon will now be part of the carbon dioxide. Likewise, all the hydrogen that was initially a part of the hydrocarbon is now a part of the water.

This means that you can use the number of moles of water and carbon dioxide, respectively, to determine how many moles of carbon and of hydrogen were originally present in the hdyrocarbon.

So for H2O you have

10.4/18.05 = 0.5761 moles

And

For CO2 you have

22.5/44 = 0.511 moles


Now you know that every mole of water contains 2mole of hydrogen and 1 mole of oxygen, which means that the reaction produced


0.5761 moles H2O *2 moles of H/1 moles of H2O

= 1.1522 moles of H


Since every mole of carbon dioxide oxide contains one mole of carbon and 2 mole of oxygen, it follows that the reaction also produced

0.511moles of CO2 *1moles C/1mole of CO2

= 0.511 moles of C

Finally to find the mole ratio that exist between C and H in the hydrocarbon divide these value by smallest one

For C = 0.511 moles/0.511moles = 1

For H = 1.1522 moles/0.511 moles = 2.25

Now for simple ratio to complete number

For C = 1*4 = 4

For H = 2.25*4 = 10.00

Hence empirical formula will be C4H10.